I have a worksheet and no book.
Points X, Y, and Z are collinear, with point X between points Y and Z. If XY=14.8 and YZ=18.2, find XZ.
I must show my work.
Will you please help.
Thanks

Answer by gonzo(654)

if they're collinear and X is between Y and Z, then the order of the points would be YXZ or ZXY depending on how you want to set them up.
i set them up as YXZ.
you're given:
XY = 14.8
YZ = 18.2
since YZ is the whole line and X is in between then YX + XZ must equal to YZ because the sum of the parts equals the whole.
substituting known values in that equation we get
YX + XZ = YZ
14.8 + XZ = 18.2
XZ = 18.2 - 14.8
XZ = 3.4
going back to the original equation and substituting for known values, we get
YX + XZ = YZ
14.8 + 3.4 = 18.2
18.2 = 18.2
XZ = 3.4 is correct.
please note YX and XY are the same line.